# Making Databases Searchable with Deep Context

### v1

@@ 1. Introduction
Fine-tuning is the other alternative to bake enterprise knowledge into the model itself. For example, Google Gemini recently topped the single-model BIRD benchmark for generating SQL queries. The idea is to curate a gold-standard dataset, run supervised fine-tuning to the train the model into a SQL specialist, and infer with self-consistency, i.e., generate several queries and select one from the group with most common result (Cloud, 2025). Lamini takes fine-tuning further by memorizing facts from the database for more effective retrieval. Unfortunately, enterprise data keeps changing, and so fine-tuning needs to be redone all the time to adapt to the latest version of the data. Furthermore, fine-tuning with either gold-standard or memorized facts is akin to caching query plans that is very hard to generalize.
@@
Fundamentally, both prompting and fine-tuning approaches “black box” natural language queries into guesswork rather than processing them systematically. This explodes the effort needed to achieve the required levels of accuracy and applicability. Furthermore, non-expert users care about the final results with SQL generation being the internal step that they cannot verify anyways. In this paper, we present Tursio, a brand new search platform to process natural language queries over databases and generate responses directly in natural language, as illustrated in Figure 1. The key idea is to connect to all databases, infer a semantic knowledge graph, contextualize user intent into the right data models using the knowledge graph, and process the queries systematically using techniques from traditional query planning and rewriting. Along the way, Tursio infuses LLMs into every part of the query processing stack, including data modeling, query compilation, query planning, and result reasoning. This has similarities to our previous work on infusing learned models into the query optimizer for better cost, cardinality, resources, and so on (Jindal et al., 2021; Wu et al., 2018; Viswanathan et al., 2018; Siddiqui et al., 2020; Negi et al., 2021).
@@ 3. Semantic Knowledge Graph
Profiling. Once a database is connected to Tursio, users can select the tables to be included in the semantic knowledge graph. We first profile the columns in each of these tables. We collect fixed-sized samples for each column, ranging from 100k to 1M rows, depending on the table sizes. Column values in production databases often have data quality issues, and so we apply a data cleaning step to gather more representative ones. Specifically, we remove outliers by filtering out the values that lie outside the inter-quartile range (IQR) for numeric values and whose length lie outside the inter-quartile range for string values. This prunes extreme low or high values, or values that are likely to be typos or invalid. We also collect statistics, e.g., counts, count distincts (full or approximate based on table sizes), min/max, average, stddev, nulls, etc.
@@ 5. Natural Language Query Planning
Tursio generates query plans from natural language queries using a systematic approach, as illustrated in Figure 4. The key idea is to break down the query planning into several steps, infusing LLMs into every step to handle the complexity of natural language. The main steps are as follows:
@@ 7. Enterprise Readiness
Maintenance. Users can train the semantic knowledge graph initially and also retrain it periodically as the underlying databases evolve. Tursio detects any manual changes to the knowledge graph and prompts users to retrain it as needed. Furthermore, we maintain a migration benchmark to maintain and migrate all prompts within the platform to newer versions of LLMs (Tripathi et al., 2025).
@@ 8.2. Table Retrieval
The results for BIRD-DEV show that Tursio achieves high overall accuracy (79%) with balanced precision (91%), recall (85%), and F1 (88%). For the BEAVER dataset family, we see a lower recall but reasonable precision, suggesting that while the model avoids retrieving irrelevant tables, additional signal (e.g., extended metadata) could further improve coverage. Still, compared to the best-performing BEAVER baseline (Precision 32%, Recall 54.4%, F1 33.9%, Perfect-Recall 15%), Tursio delivers substantial gains: approximately 55% improvement in precision, along with clear improvements in both F1 and Perfect-Recall.
@@ 8.3. Join Inference
For BIRD-DEV and TPC-H, the ground truth consists of the full set of joins defined in the official schema metadata, including both explicit primary–foreign key pairs and additional schema-declared referential constraints. For the BEAVER benchmark, we use all annotated join pairs provided for both warehouses. For the internal production workload, join annotations were curated by domain experts. During evaluation, we compare these ground-truth pairs with the join connections inferred by our semantic knowledge-graph construction. We report the resulting precision, recall, and F1 scores in Figures 9,  10, and 11. The BEAVER schemas represent the most challenging setting, exhibiting limited primary–foreign key discipline. Consequently, our model achieves low average Accuracy (14.7%), Precision (40%), and Recall (17.3%), as it struggles to distinguish true join paths from semantically plausible alternatives in loosely structured schemas. In contrast, performance on public benchmark datasets is substantially stronger. On BIRD, the model attains an average Accuracy of 50.46%, Precision of 82.44%, and Recall of 58.93%. These improvements stem from the clearer relational structure and more consistent key constraints, which enable recovery of most valid join relationships with high precision. Nonetheless, imperfect normalization in some cases prevents us from having a better recall.
TPC-H and production workloads further illustrate the model’s upper-bound performance in well-engineered environments. These schemas are normalized and stable, allowing the model to achieve its strongest results—Accuracy exceeding 63%, Precision reaching 100%, and Recall between 67% and 78%.
@@ 8.6. Auxiliary Capabilities
Summary Quality. To assess the quality of automatically generated summaries, we employ an LLM-as-a-grader evaluation protocol using GPT-4.1 and GPT-5. The grader is provided with a structured rubric that categorizes each summary into one of four quality levels: Excellent, Good, Average, or Poor. The rubric evaluates five dimensions—clarity and coherence, conciseness, structure, factuality, and length appropriateness—and requires the model to output both a categorical rating and a short justification. Each categorical label is mapped to a numerical score using a fixed scale (Excellent = 1, Good = 2, Average = 3, Poor = 4). For every summary, both models independently produce a score, and we report the average of their individual scores.
The summary quality evaluation, via dual LLM grading, achieves an average score of 2.2, suggesting high-quality structured summaries.
PII Detection. We compute the precision and recall over labeled PII entities to evaluate whether the system successfully detects sensitive columns while minimizing false positives. Our results show that PII detection achieves strong recall (95%) and precision (91%), which is suitable for production-grade privacy constraints.
@@ 8.7. Case Studies
Case study 1: Serving customer tickets. A large home warranty company wants to improve their customer experience with better ticket handling. Their support platform generates templated queries that must fetch relevant data, analyze the typical resolution, apply business rules to ensure the promised customer agreements are respected, and finally generate a response to resolve the request. For each response generated by Tursio, a human agent verifies its correctness (marking it as helpful or not helpful) before sending it to the customer. Table 2 shows the human-evaluated accuracy over a recent six-week period of production usage. Serving hundreds of queries per week, Tursio achieves over 96% raw accuracy and 100% effective accuracy (excluding cases when insufficient information was provided as input) consistently.
Case study 2: Dynamic dashboard querying. A supply chain and planning company wants to allow its users to dynamically query their data along various planning points. Their existing dashboards are static and do not allow users to ask ad-hoc questions, and they require natural language queries to be interpreted based on their underlying data model. They ran an exhaustive validation using a set of 192 queries crafted by their domain experts over their data model. Table 3 shows the accuracy of Tursio in terms of compilation and validation over these queries. Tursio achieves 86% compilation accuracy, which goes up to 88% when excluding cases with missing data. Furthermore, when validating queries while excluding those with missing data or metrics, Tursio achieves 94% validation accuracy.
@@
9. The Road Ahead
@@
Search for all data. Modern users don’t care where the data lives, they just want to search it and get answers. Tursio currently connects to multiple databases and data warehouses, but future work will focus on connecting to unstructured and even web data. Identifying possible way to link structured and unstructured data together will be a key research challenge.

### v2

@@ 1. Introduction
Fine-tuning is the other alternative to improve SQL generation accuracy. For example, Google Gemini recently topped the single-model BIRD benchmark by curating gold-standard datasets and running supervised fine-tuning to produce a stronger SQL specialist, combined with self-consistency inference (Cloud, 2025). While such general-purpose fine-tuning improves the model’s SQL fluency broadly, it does not encode the enterprise-specific context—proprietary schemas, business logic, domain terminology—that is essential for accurate query generation over internal databases. Lamini takes a different approach by memorizing facts from specific databases for more targeted retrieval. However, enterprise data keeps changing, and memorized facts need to be refreshed continually. More fundamentally, memorizing facts or query patterns does not compose: the number of possible query plans grows combinatorially with schema size, making exhaustive memorization impractical.
@@
Fundamentally, both prompting and fine-tuning “black box” natural language queries into guesswork rather than processing them systematically. This explodes the effort needed to achieve the required levels of accuracy and applicability. Furthermore, non-expert users care about the final results with SQL generation being the internal step that they cannot verify anyways. In this paper, we present Tursio, a brand new search platform to process natural language queries over databases and generate responses directly in natural language, as illustrated in Figure 1. The key idea is to connect to all databases, infer a semantic knowledge graph, contextualize user intent into the right data models using the knowledge graph, and process the queries systematically using techniques from traditional query planning and rewriting. Along the way, Tursio infuses LLMs into every part of the query processing stack, including data modeling, query compilation, query planning, and result reasoning. Philosophically, this mirrors the idea from our prior work on learned query optimization (Jindal et al., 2021; Wu et al., 2018; Negi et al., 2021): rather than relying on a single monolithic model, decomposing the problem into fine-grained steps—each augmented with a targeted model—yields better accuracy and debuggability.
@@ 3. Semantic Knowledge Graph
Profiling. Once a database is connected to Tursio, users can select the tables to be included in the semantic knowledge graph. We first profile the columns in each of these tables. We collect fixed-sized samples for each column, ranging from 100k to 1M rows, depending on the table sizes. Column values in production databases often have data quality issues, and so we apply a data cleaning step to gather more representative ones. Specifically, we remove outliers by filtering out the values that lie outside the inter-quartile range (IQR) for numeric values and whose length lie outside the inter-quartile range for string values. This prunes extreme low or high values, or values that are likely to be typos or invalid. We note that profiling focuses on representative value distributions and may not capture all data quality patterns, e.g., when the same information exists in both raw and cleaned columns. However, the column descriptions and LLM-based reasoning in subsequent steps help disambiguate such cases by comparing column semantics beyond their value distributions alone. We also collect statistics, e.g., counts, count distincts (full or approximate based on table sizes), min/max, average, stddev, nulls, etc.
@@ 5. Natural Language Query Planning
Tursio generates query plans from natural language queries using a systematic approach, as illustrated in Figure 4. The key idea is to break down the query planning into several steps, infusing LLMs into every step to handle the complexity of natural language.
A natural question is why not simply prompt an LLM with the relevant schema context and let it generate SQL directly. While end-to-end approaches work for simpler queries, they suffer from compounding errors on complex enterprise schemas: the model must simultaneously resolve table selection, join paths, column mapping, filter extraction, and aggregation semantics in a single generation step, with no opportunity to catch intermediate mistakes. Tursio’s multi-step approach constrains the problem at each stage: parsed operators are validated by deterministic ANTLR parsers, operands are grounded against the knowledge graph metadata, and the operator tree enforces relational algebra rules. Each step narrows the space of valid outputs for the next, making errors detectable and correctable rather than silently propagated into the final SQL. This decomposition also enables the query breakdown feature (Section 6), where users can see exactly which parts of their question were interpreted and how.
The main steps are as follows:
@@ 7. Enterprise Readiness
Maintenance. Users can train the semantic knowledge graph initially and also retrain it periodically as the underlying databases evolve. Tursio detects any manual changes to the knowledge graph and prompts users to retrain it as needed. To manage the cost of retraining, Tursio supports incremental updates: when schemas change, only the affected tables and their relationships are re-profiled and re-inferred, avoiding the need for full retraining. Tursio tracks schema versions and computes deltas to determine which portions of the knowledge graph require updates, significantly reducing the number of LLM calls for routine schema evolution. Furthermore, we maintain a migration benchmark to maintain and migrate all prompts within the platform to newer versions of LLMs (Tripathi et al., 2025).
@@ 8.2. Table Retrieval
The results for BIRD-DEV show that Tursio achieves high overall accuracy ($85\%$) with balanced precision ($94\%$), recall ($89\%$), F1 ($92\%$), and Perfect-Recall ($81\%$). For the BEAVER dataset family, we see a lower recall but reasonable precision, suggesting that while the model avoids retrieving irrelevant tables, additional signal (e.g., extended metadata) could further improve coverage. Still, compared to the best-performing BEAVER baseline (Precision $32\%$, Recall $54.4\%$, F1 $33.9\%$, Perfect-Recall $15\%$), Tursio delivers substantial gains: in Precision and F1 scores ($91\%$ and $50\%$ respectively), marginal improvement in Perfect-Recall to $15.33\%$. The lower recall of $38\%$ indicates that our model is more conservative in retrieving tables, which may be a trade-off to maintain high precision in the face of noisy or ambiguous schemas. Here we compare against best-performing baselines for BEAVER which include the best of all the models evaluated in the BEAVER paper, across both the top-10 and top-5 retrieval performance. In our table retrieval algorithm, we retrieve the exact tables that are required to answer the query, which is a more stringent requirement than retrieving a superset of relevant tables. This leads to higher precision but can reduce recall if the model fails to identify all necessary tables. Overall, the results demonstrate that Tursio achieves strong table retrieval performance across both the benchmark datasets.
@@ 8.3. Join Inference
For BIRD-DEV and TPC-H, the ground truth consists of the full set of joins defined in the official schema metadata, including both explicit primary–foreign key pairs and additional schema-declared referential constraints. For the BEAVER benchmark, we use all annotated join pairs provided for both warehouses. For the internal production workload, join annotations were curated by domain experts. During evaluation, we compare these ground-truth pairs with the join connections inferred by our semantic knowledge-graph construction. We report the resulting precision, recall, and F1 scores in Figures 9,  10, and 11. The BEAVER schemas represent the most challenging setting, exhibiting limited primary–foreign key discipline. Consequently, our model achieves low average Accuracy ($14.7\%$), Precision ($40\%$), and Recall ($17.3\%$), as it struggles to distinguish true join paths from semantically plausible alternatives in loosely structured schemas. In contrast, performance on public benchmark datasets is substantially stronger. On BIRD, the model attains an average Accuracy of $50.46\%$, Precision of $82.44\%$, and Recall of $58.93\%$. These improvements stem from the clearer relational structure and more consistent key constraints, which enable recovery of most valid join relationships with high precision. Nonetheless, imperfect normalization in some cases prevents us from having a better recall.
TPC-H and production workloads further illustrate the model’s upper-bound performance in well-engineered environments. These schemas are normalized and stable, allowing the model to achieve its strongest results—Accuracy exceeding $63\%$, Precision reaching $100\%$, and Recall between $67\%$ and $78\%$.
@@ 8.6. Auxiliary Capabilities
Summary Quality. To assess the quality of automatically generated summaries, we employ an LLM-as-a-grader evaluation protocol using GPT-4.1 and GPT-5. The grader is provided with a structured rubric that categorizes each summary into one of four quality levels: Excellent, Good, Average, or Poor. The rubric evaluates five dimensions—clarity and coherence, conciseness, structure, factuality, and length appropriateness—and requires the model to output both a categorical rating and a short justification. Each categorical label is mapped to a numerical score using a fixed scale (Excellent = 1, Good = 2, Average = 3, Poor = 4). For every summary, both models independently produce a score, and we report the average of their individual scores. The summary quality evaluation, via dual LLM grading, achieves an average score of 2.2, suggesting high-quality structured summaries.
PII Detection. We compute the precision and recall over labeled PII entities to evaluate whether the system successfully detects sensitive columns while minimizing false positives. Our results show that PII detection achieves strong recall ($95\%$) and precision ($91\%$), which is suitable for production-grade privacy constraints.
@@ 8.7. Case Studies
Case study 1: Serving customer tickets. A large home warranty company wants to improve their customer experience with better ticket handling. Their support platform generates templated queries that must fetch relevant data, analyze the typical resolution, apply business rules to ensure the promised customer agreements are respected, and finally generate a response to resolve the request. The underlying schema spans over several tables covering warranty contracts, service history, customer profiles, and coverage terms, with domain-specific terminology (e.g., “covered items”, “courtesy amount”) that general-purpose LLMs would not understand out of the box. Tursio’s semantic knowledge graph automatically mapped these domain terms to the correct columns and inferred the join paths between contract, service, and customer tables. Custom measures were defined to encode business rules such as coverage eligibility and service fee calculations, ensuring that the generated responses respect the contractual agreements. For each response generated by Tursio, a human agent verifies its correctness (marking it as helpful or not helpful) before sending it to the customer. Table 2 shows the human-evaluated accuracy over a recent six-week period of production usage. Serving hundreds of queries per week, Tursio achieves over $96\%$ raw accuracy and $100\%$ effective accuracy (excluding cases when insufficient information was provided as input) consistently. The gap between raw and effective accuracy is attributed entirely to incomplete input tickets that lack required details (e.g., missing diagnostic details), rather than errors in query planning or result generation.
Case study 2: Dynamic dashboard querying. A supply chain and planning company wants to allow its users to dynamically query their data along various planning points. Their existing dashboards are static and do not allow users to ask ad-hoc questions, and they require natural language queries to be interpreted based on their underlying data model. The data model contains wide planning tables (50+ columns) with complex hierarchical dimensions (product categories, geographic regions, time periods) and domain-specific metrics. Tursio’s join inference was critical for connecting the planning dimension tables, while auto-generated sample questions provided grounding for the diverse ways domain experts phrase analytical queries over these dimensions. They ran an exhaustive validation using a set of 192 queries crafted by their domain experts over their data model. Table 3 shows the accuracy of Tursio in terms of compilation and validation over these queries. Tursio achieves $86\%$ compilation accuracy, which goes up to $88\%$ when excluding cases with missing data. Furthermore, when validating queries while excluding those with missing data or metrics, Tursio achieves $94\%$ validation accuracy. An analysis of the failure scenarios revealed that the majority involved metrics not yet modeled in the knowledge graph or queries requiring multi-step reasoning across temporal hierarchies—areas targeted for improvement in future iterations.
@@
9. Related Work & Road Ahead
Text-to-SQL. Translating natural language to SQL has been extensively studied, with benchmarks like Spider (Yu et al., 2018) and BIRD (Li et al., 2023) driving rapid progress. Recent LLM-based approaches include DIN-SQL (Pourreza and Rafiei, 2023), which decomposes the task into sub-problems, DAIL-SQL (Gao et al., 2024), which optimizes example selection for in-context learning, CHESS (Talaei et al., 2024), which adds contextual retrieval and schema pruning, and MAC-SQL (Wang et al., 2024), which employs a multi-agent architecture. While these approaches focus on generating SQL from natural language, they assume a given schema and do not address the upstream challenge of automatically inferring semantic context from the raw database. Tursio targets the broader end-to-end problem: from automatic semantic modeling to query planning to delivering natural language answers rather than SQL. Furthermore, Tursio’s multi-step planning with deterministic grounding constrains LLM outputs at each stage, unlike end-to-end approaches that rely solely on the model to produce valid SQL in a single generation step.
Data Discovery. Discovering relationships across tables is central to data integration and search. Aurum (Fernandez et al., 2018) builds an enterprise knowledge graph for data discovery using profiling signals. Valentine (Koutras et al., 2021) provides a comprehensive benchmark for schema matching techniques. Starmie (Fan et al., 2023) uses pre-trained language models for column-level data discovery in data lakes. Tursio’s join inference builds on inclusion dependency discovery (Rostin et al., 2009) and extends it with LLM-based validation, combining statistical signals (approximate count distinct overlap) and semantic reasoning (LLM as judge) for robust primary-key/foreign-key detection. This hybrid approach allows Tursio to scale join inference to larger schemas while maintaining high precision.
Semantic Layers and NL Interfaces. Semantic layers such as dbt Semantic Layer, LookML, and Cube provide business-friendly abstractions over databases but require manual modeling and ongoing maintenance. Tursio automates this step using LLM-powered inference over the raw database schema and data profiles. Earlier NL database interfaces like NaLIR (Li and Jagadish, 2014) used interactive query tree disambiguation over dependency parses, while neural semantic parsing approaches like RAT-SQL (Wang et al., 2020) pioneered relation-aware schema encoding and linking. Tursio combines ideas from both traditions: it uses schema-aware encoding via the semantic knowledge graph (similar to neural approaches) while preserving the ability to decompose and validate query plans step-by-step (similar to rule-based approaches), resulting in a system that is both accurate and explainable.
@@ 9. Related Work & Road Ahead
Search for all data. Modern users don’t care where the data lives, they just want to search it and get answers. While Tursio currently supports both SQL and NoSQL data stores, our evaluation has primarily focused on relational databases. Extending the evaluation to NoSQL systems (e.g., Cassandra, CosmosDB) and systematically assessing how query planning generalizes across database dialects is an important next step. Beyond existing connectors, future work will focus on connecting to unstructured and even web data. Identifying possible ways to link structured and unstructured data together will be a key research challenge.